what is the range of y=3x+2 if the domain is {-3,-1,0,1,3}?
Answer by arodr172(1)

To solve this, you have to plug-in each value of the domain to see what the y values are. The result of all your y values is the set of the range.
Equation:
y = 3x + 2 if the domain is {-3,-1,0,1,3}
Solution:
y = 3(-3) + 2 => -9 + 2 = -7
y = 3(-1) + 2 => -3 + 2 = -1
y = 3(0) + 0 = 0
y = 3(1) + 2 = 5
y = 3(3) + 2 = 11
Therefore, the range is the set of all the y values {-7, -1, 0, 5, 11}.